# Evaluating Quantumness, Efficiency and Cost of Quantum Random Number Generators via Photon Statistics

### v5

@@ VIII Performance and Cost of the QRNG
Considering all the discussions in the previous section, we can now write random number generation rate $r$ as a function of randomness. From Eq. (31), (36) and (47), we have
@@
[…]
@@
$H_{min}=\log\frac{8\epsilon}{(1+2\epsilon)^{2}\left(1-\left(\frac{1-2\epsilon}{1+2\epsilon}\right)^{\frac{2}{N}}\right)}$ (54)
@@
$\mu$uparrow $\downarrow$ $\uparrow$
$T$uparrow -- --
$d$uparrow $\downarrow$ $\downarrow$
@@
The above discussion has been described as a heat map in Fig. 14. In the figure, dark shade denotes low value and light shade denotes high value of $\epsilon$, minimum entropy, generation rate and cost in Figures 14(a), 14(b), 14(c) and 14(d), respectively. Figure 14(a) and (b) shows that high randomness demands a small value of the product $\mu Td$. However, the product $\mu d$ cannot be very small as Fig. 14(b) indicates. A small $\mu d$ would result in a low random number generation rate. On the other hand, to keep the cost of the QRNG low, we need a small $d$ but a large $\mu$.
@@
IX Results of Randomness Tests
Figure 15: NIST result. $p$-values for passed tests are plotted as bar charts for (a) $\mu Td=0.05$, (b) $\mu Td=0.10$, (c) $\mu Td=0.15$, (d) $\mu Td=0.20$, (e) $\mu Td=0.25$, (f) $\mu Td=0.30$, (g) $\mu Td=0.35$, (h) $\mu Td=0.40$ , (i) $\mu Td=0.45$, (j) $\mu Td=0.50$. For each $\mu Td$, three $1$ GB files are generated by simulating probability distribution (31), and the NIST tests have been performed on those files. Different textures denote results corresponding to different files. A test is considered as passed if the corresponding $p$-value is at least $0.01$. Tests that fail for all of the three files are not included in the bar chart. The vertical dashed line denotes $p$-value $=0.01$
@@ IX Results of Randomness Tests
Figure 16: NIST test results for the data simulated with the distribution given by Eq. (31). Three files, each containing $1$ GB random number have been used for the test for every $\mu Td$. A test fails if the corresponding $p$-value is found less than $0.01$. The range in the second column arises due to different results from three different files. The number of passed tests decreases with the increment of $\mu Td$. This indicates that as $\mu Td$ increases, the randomness in the generated random number decreases.
The NIST Randomness Testing Suite is an extensive set of statistical tests developed by the National Institute of Standards and Technology (NIST) to evaluate the quality of randomness in random numbers. It is one of the most widely used suites for testing randomness. It contains $15$ different tests to examine different aspects of randomness. Most of the tests have the standard normal or $\chi^{2}$ as reference distribution for their test statistic. Here null hypothesis is taken as the number to be tested is random. For each test, the significance level is considered as $0.01$. If the value of the test statistic for some tests falls in the critical region, or equivalently, calculated $p$-value is less than $0.01$, then the corresponding tests fail rejecting the null hypothesis. A perfectly random number should pass all of the $15$ tests. As mentioned in Section V, the central limit theorem is applicable only for large $n$, tests that use the central limit theorem to determine the reference distribution, have a requirement for the bit stream, that is, the length of the testing sequence, to be large enough (at least $10^{6}$).
@@
Figure 17: (a) Dieharder and (b) AIS-$31$ test results for the data simulated with the distribution given by Eq. (31). For Dieharder, there are $30$ tests; for AIS-$31$, $9$ tests are there. These tests also reveal that the number of passed tests decreases as $\mu Td$ increases indicating lower values of $\mu Td$ result in better randomness before post-processing. Along with pass and fail, Dieharder tests also mention suspected results as weak. They are also included in this figure.
We have generated random numbers simulating the probability distribution given by Eq. (31) for different $\mu Td$. Three files of size $1$ GB have been generated for every $\mu Td$. Then we perform the NIST tests over the generated numbers. $100$ bit streams of length $10,000,000$ have been considered for the tests. In Fig. 15, we plot the bar charts of the $p$-values from different NIST tests. The tests for which the $p$-values are found as 00 are not included in the plot. From these plots, we can see that, as $\mu Td$ increases, the number of passed tests decreases indicating a decrement of the randomness in the random number. The overall result is plotted in Fig. 16 as the number of tests our simulated data passed. These graphs indicate that lower $\mu Td$ values give rise to better random numbers requiring less post-processing.
@@ X Discussion
Appendix A A Comparison between trusted QRNG and DI-QRNG
A device-independent quantum protocol uses quantum entanglement to ensure security. This can be verified using Bell inequality, quantum discord, etc. acin2016certified ; Alicki_2008 ; Alicki_2008a ; Martin_2023 . On the other hand, if a protocol relies on single photons as a resource, the absence of entanglement prevents the application of the validation procedures mentioned above. However, we can still verify its quantumness using photon-count statistics PhysRevA.41.475 ; fox2006quantum ; RevModPhys.68.127 ; qute.202100062 , discussed in Section  II and B, to improve the security. A detailed comparison indicating advantages and disadvantages is given in Table 2.
@@ Appendix A A Comparison between trusted QRNG and DI-QRNG
Appendix B Relationship between Non-classicality of Light, antibunching of Photons and sub-Poissonian Distributions
The second-order correlation function fox2006quantum of light is defined by
[…]
and $\text{and }g^{2}(\tau)$ $=1\text{ for }\tau\gg\tau_{c},$
$\tau_{c}$ being the coherence time of the source.
where $n_{i}(t)$ is the number of counts registered on detector $i$ at time $t$. If the incoming light consists of a photon stream with long intervals between two consecutive photons, the product $n_{1}(t)n_{2}(t)$ is always 00, giving $g^{2}(0)=0<1$, which is not possible with classical lights. Thus, depending on $g^{2}(t)$, light is distinguished in the following three classes fox2006quantum :
  1. (i)
    bunched light, when $g^{2}(0)>1$,
  2. (ii)
    coherent light, when $g^{2}(0)=1$,
  3. (iii)
    antibunched light, when $g^{2}(0)<1$.
Out of these three, antibunched light is undoubtedly non-classical.
[…]
$g^{2}(\tau)=\frac{1}{2}\cos(\omega_{1}-\omega_{2})\tau-\frac{1}{n}+1,$ (59)
@@ Appendix C Statistical Tools
C.1 Interval Estimation
In statistics, an interval $(a,b)$ is called the confidence interval for a distribution parameter $\Theta$, if $P(A<\Theta<B)=1-\epsilon$ for some predefined $\epsilon\in(0,1)$, where $A,B$ are the random variables corresponding to the statistics $a,b$ (they are functions of sample values) respectively. The quantity $1-\epsilon$ is called confidence level of the interval. The procedure of finding the confidence interval for the parameter $\Theta$ is called interval estimation. A statistic $a$ is called an unbiased estimate of an unknown parameter $\Theta$ if the expectation $\langle A\rangle=\Theta$. A statistic $a$ is called a consistent estimate of a parameter $\Theta$ if for every real $\epsilon>0$, $\lim\limits_{n\to\infty}P(|A-\Theta|\geq\epsilon)=0$, where $n$ is the sample size.
C.2 Hypothesis Testing
In statistics, a hypothesis is a statement or an assumption about a population distribution. this may be right or wrong. Hypothesis testing is a procedure to decide whether a statistical hypothesis, called the null hypothesis $H_{0}$, should be rejected or not. A critical region or rejection region of hypothesis testing is a specific region of test statistic values, for which the null hypothesis is rejected. The test statistic for a hypothesis test is a function of the testing sample casella2002statistical . A hypothesis test is called one-tailed if the critical region appears only at one tail of the reference distribution of the test statistic. On the other hand, it is called a two-tailed test, if the critical regions appear at both tails of the above distribution.
During hypothesis testing, two types of error may occur. Type I error occurs when $H_{0}$ is rejected given that it is true. Type II error occurs when $H_{0}$ is accepted given that it is false. Let $\alpha$ and $\beta$ be the probabilities of Type I and Type II errors respectively. The value $\alpha$ denotes the area of the critical region and is also called the significance level. The value $1-\beta$, that is, the probability that $H_{0}$ is rejected given that it is false, is also called the power of a test.
@@ Appendix E Randomness Testing Suites
To overcome this limitation, several randomness testing suites like NIST 10.5555/2206233 , Dieharder Die , AIS-$31$ killmann2011proposal , ENT ENT etc. have been proposed. Recently, Foreman et al. also suggested some test suite foreman2024 for testing randomness. Each of these suits consists of multiple tests to decide the randomness of the input data. For each test, they calculate the value of the test statistic, and depending on these values the randomness is decided. NIST and Dieharder perform hypothesis testing with the null hypothesis that the number is random, AIS-31 performs interval estimation, and ENT directly returns the test statistic values. These suites consider different aspects, such as entropy, serial correlation, frequency of values, repetition of similar patterns etc. to perform the test. Some tests also perform Monte Carlo simulation MonteCarlo to validate the randomness.

### v6

@@ VI Simultaneous Analysis of Rate, Randomness, Entropy and Cost
Considering all the discussions in the previous section, we can now write random number generation rate $r$ as a function of randomness. From Eq. (22), (26) and (37), we have,
@@
$r_{max}=\frac{2\log N}{Nk_{tol}\delta_{t}}\ln{\frac{1+2\epsilon}{1-2\epsilon}}\text{ bits/ unit time},$ (43)
@@
$H_{min}=\log\frac{8\epsilon}{(1+2\epsilon)^{2}\left(1-\left(\frac{1-2\epsilon}{1+2\epsilon}\right)^{\frac{2}{N}}\right)}.$ (44)
@@
$\mu$ $\uparrow$ $\downarrow$ $\uparrow$
$T$ $\uparrow$ - -
$d$ $\uparrow$ $\downarrow$ $\downarrow$
@@
The above discussion has been described as a heat map in Fig. 13. In the figure, dark shade denotes low value and light shade denotes high value of $\epsilon$, minimum entropy, generation rate and cost in Figures 13(a), 13(b), 13(c) and 13(d), respectively. Figure 13(a) and (b) shows that high randomness demands a small value of the product $\mu Td$. However, the product $\mu d$ cannot be very small as Fig. 13(b) indicates. A small $\mu d$ would result in a low random number generation rate. On the other hand, to keep the cost of the QRNG low, we need a small $d$ but a large $\mu$.
@@
VII Results of Randomness Tests
Figure 14: NIST result. $p$-values for passed tests are plotted as bar charts for (a) $\mu Td=0.05$, (b) $\mu Td=0.10$, (c) $\mu Td=0.15$, (d) $\mu Td=0.20$, (e) $\mu Td=0.25$, (f) $\mu Td=0.30$, (g) $\mu Td=0.35$, (h) $\mu Td=0.40$. For each $\mu Td$, three 11 GB files are generated by simulating probability distribution (22), and the NIST tests have been performed on those files. Different textures denote results corresponding to different files. A test is considered as passed if the corresponding $p$-value is at least $0.01$. Tests that fail for all of the three files are not included in the bar chart. The vertical dashed line denotes $p$-value =0.01=0.01
@@ VII Results of Randomness Tests
Figure 15: NIST test results for the data simulated with the distribution given by Eq. (22). Three files, each containing 11 GB random number have been used for the test for every $\mu Td$. A test fails if the corresponding $p$-value is found less than $0.01$. The range in the second column arises due to different results from three different files. The number of passed tests decreases with the increment of $\mu Td$. This indicates that as $\mu Td$ increases, the randomness in the generated random number decreases.
The NIST Randomness Testing Suite is an extensive set of statistical tests developed by the National Institute of Standards and Technology (NIST) to evaluate the quality of randomness in random numbers. It is one of the most widely used suites for testing randomness. It contains $15$ different tests to examine different aspects of randomness. Most of the tests have the standard normal or $\chi^{2}$ as reference distribution for their test statistic. Here null hypothesis is taken as the number to be tested is random. For each test, the significance level is considered as $0.01$. If the value of the test statistic for some tests falls in the critical region, or equivalently, calculated $p$-value is less than $0.01$, then the corresponding tests fail rejecting the null hypothesis. A perfectly random number should pass all of the $15$ tests. As mentioned in Section III, the central limit theorem is applicable only for large $n$, tests that use the central limit theorem to determine the reference distribution, have a requirement for the bit stream, that is, the length of the testing sequence, to be large enough (at least $10^{6}$).
Figure 16: (a) Dieharder and (b) AIS-$31$ test results for the data simulated with the distribution given by Eq. (22). For Dieharder, there are $30$ tests; for AIS-$31$, 99 tests are there. These tests also reveal that the number of passed tests decreases as $\mu Td$ increases, indicating that lower values of $\mu Td$ result in better randomness before post-processing. Along with pass and fail, Dieharder tests also mention suspected results as weak. They are also included in this figure.
@@
We have generated random numbers simulating the probability distribution given by Eq. (22) for different $\mu Td$. Three files of size $1$ GB have been generated for every $\mu Td$. Then we perform the NIST tests over the generated numbers. $100$ bit streams of length $10,000,000$ have been considered for the tests. In Fig. 14, we plot the bar charts of the $p$-values from different NIST tests. The tests for which the $p$-values are found as 0 are not included in the plot. From these plots, we can see that, as $\mu Td$ increases, the number of passed tests decreases, indicating a decrement of the randomness in the random number. The overall result is plotted in Fig. 15 as the number of tests our simulated data passed. These graphs indicate that lower $\mu Td$ values give rise to better random numbers requiring less post-processing.
@@ VIII Discussion
Acknowledgement
S. Das would like to acknowledge the support received by the Quantum Delta NL KAT-2 project.
Data Availability
Datasets that have been used to validate the two-fold method and the Python codes which are used to generate these datasets are available at
https://github.com/sp-k/QRNG_eval.git.
Appendix A Relationship between Non-classicality of Light, antibunching of Photons and sub-Poissonian Distributions
The distribution of the photon-counting statistics can be divided into three classes fox2006quantum :
  1. (i)
  2. (ii)
  3. (iii)
Classical wave theory of light follows Poissonian and super-Poissonian statistics fox2006quantum . On the other hand, the sub-Poissonian nature of a photon source cannot be explained by the classical wave theory, making it a clear indication of its quantum nature RevModPhys.68.127 ; qute.202100062 .
For example, thermal light follows Bose–Einstein distribution fox2006quantum , given by
$P(n)=\frac{1}{\mu+1}\left(\frac{\mu}{\mu+1}\right)^{n}.$ (45)
$P(n,T)=\frac{e^{-\mu T}(\mu T)^{n}}{n!},$ (47)
Antibunching property can also be used to distinguish quantum light from classical light. The second-order correlation function fox2006quantum of light is defined by
where $I(t)$ is the intensity of the light at time $t$ and $\langle\cdot\rangle$ denotes the average over all $t$. Using the classical theory of light, it can be shown that RevModPhys.68.127
$\text{and }g^{2}(\tau)$ $=1\text{ for }\tau\gg\tau_{c},$
$\tau_{c}$ being the coherence time of the source.
where $n_{i}(t)$ is the number of counts registered on detector $i$ at time $t$. If the incoming light consists of a photon stream with long intervals between two consecutive photons, the product $n_{1}(t)n_{2}(t)$ is always 0, giving $g^{2}(0)=0<1$, which is not possible with classical lights. Thus, depending on $g^{2}(t)$, light is distinguished in the following three classes fox2006quantum :
  1. (i)
    bunched light, when $g^{2}(0)>1$,
  2. (ii)
    coherent light, when $g^{2}(0)=1$,
  3. (iii)
    antibunched light, when $g^{2}(0)<1$.
Out of these three, antibunched light is undoubtedly non-classical.
$\langle(\Delta N)^{2}\rangle-\langle N\rangle=\langle N\rangle^{2}\frac{1}{T^{2}}\int_{-T}^{T}d\tau(T-|\tau|)(g^{2}(\tau)-1).$ (51)
$g^{2}(\tau)=\frac{1}{2}\cos(\omega_{1}-\omega_{2})\tau-\frac{1}{n}+1,$ (52)
Appendix B A Comparison between trusted QRNG and DI-QRNG
@@
A device-independent quantum protocol uses quantum entanglement to ensure security. This can be verified using Bell inequality, quantum discord, etc. acin2016certified ; Alicki_2008 ; Alicki_2008a ; Martin_2023 . On the other hand, if a protocol relies on single photons as a resource, the absence of entanglement prevents the application of the validation procedures mentioned above. However, we can still verify its quantumness using photon-count statistics PhysRevA.41.475 ; fox2006quantum ; RevModPhys.68.127 ; qute.202100062 , discussed in A, to improve the security. A detailed comparison indicating advantages and disadvantages is given in Table 3.
@@ Appendix C Statistical Tools
C.1 Point Ectimation
In statistics, a statistic $a$, which is a function of sample values, is called an estimator for a distribution parameter $\Theta$, if $P(A=\Theta)=1-\epsilon$ for some given $\epsilon\in(0,1)$, where $A$ is the random variable corresponding to the statistic $a$. The procedure of finding such an estimator is called point estimation. The quantity $1-\epsilon$ is called confidence level of the estimator. A statistic $a$ is called an unbiased estimate of an unknown parameter $\Theta$ if the expectation $\langle A\rangle=\Theta$. A statistic $a$ is called a consistent estimate of a parameter $\Theta$ if for every real $\epsilon>0$, $\lim\limits_{n\to\infty}P(|A-\Theta|\geq\epsilon)=0$, where $n$ is the sample size.
C.2 Interval Estimation
In statistics, an interval $(a,b)$ is called the confidence interval for a distribution parameter $\Theta$, if $P(A<\Theta<B)=1-\epsilon$ for some predefined $\epsilon\in(0,1)$, where $A,B$ are the random variables corresponding to the statistics $a,b$ (they are functions of sample values) respectively. The quantity $1-\epsilon$ is called confidence level of the interval. The procedure of finding the confidence interval for the parameter $\Theta$ is called interval estimation.
C.3 Hypothesis Testing
In statistics, a hypothesis is a statement or an assumption about a population distribution, this may be right or wrong. Hypothesis testing is a procedure to decide whether a statistical hypothesis, called the null hypothesis $H_{0}$, should be rejected or not. A critical region or rejection region of hypothesis testing is a specific region of test statistic values, for which the null hypothesis is rejected. The test statistic for a hypothesis test is a function of the testing sample casella2002statistical . A hypothesis test is called one-tailed if the critical region appears only at one tail of the reference distribution of the test statistic. On the other hand, it is called a two-tailed test, if the critical regions appear at both tails of the above distribution.
During hypothesis testing, two types of error may occur. Type I error occurs when $H_{0}$ is rejected given that it is true. Type II error occurs when $H_{0}$ is accepted given that it is false. Let $\alpha$ and $\beta$ be the probabilities of Type I and Type II errors, respectively. The value $\alpha$ denotes the area of the critical region and is also called the significance level. The value $1-\beta$, that is, the probability that $H_{0}$ is rejected given that it is false, is also called the power of a test.
@@ Appendix E Randomness Testing Suites
To overcome this limitation, several randomness testing suites like NIST 10.5555/2206233 , Dieharder Die , AIS-$31$ killmann2011proposal , ENT ENT etc. have been proposed. Recently, Foreman et al. also suggested some test suite foreman2024 for testing randomness. Each of these suits consists of multiple tests to decide the randomness of the input data. For each test, they calculate the value of the test statistic, and depending on these values the randomness is decided. NIST and Dieharder perform hypothesis testing with the null hypothesis that the number is random, AIS-31 performs interval estimation, and ENT directly returns the test statistic values. These suites consider different aspects, such as entropy, serial correlation, frequency of values, repetition of similar patterns etc., to perform the test. Some tests also perform Monte Carlo simulation MonteCarlo to validate the randomness.